INFRARED PHOTODISSOCIATION SPECTROSCOPY OF PROTONATED METHANOL CLUSTER IONS

KENSUKE TONO and \underline{KOICHI TSUKIYAMA}, Department of Chemistry, Faculty of Science, Tokyo University of Science, 1-3 Kagurazaka, Shinjuku, Tokyo, 162-8601, Japan.

Photodissociation spectra of the protonated methanol cluster ions, H^{+}(CH_{3}OH)_{n} (n = 4−8), were measured in the range of 900 − 2300 cm^{−1} by using a infrared free electron laser. A broad spectral feature was observed at about 1600 cm^{−1} in each spectrum. We assign this feature to the OH^{+} stretching vibration, which is coupled with the OH bending motion in the HOH^{+} group. The small vibrational wavenumber suggests that the O-H^{+} bond is very weak. Structure optimizations by the density functional theory show that H^{+} is equally shared by two oxygen atoms at n = 4, 5. This result also demonstrates the weak bonding between H^{+} and oxygen. The binding features of H^{+} in the cluster ions will be discussed.